0010 UT   Trying to wake MODS, but it hands during 'blue tedpower on'
0020 UT   Attempting to reset MODS to clear the problem w/ 'tedpower'
0030 UT   Test preset to TCS
0058 UT   mods1b.20130412.0003-5 have a large artifact at the blue ends of the spectra due to light leaks from solar illumination inside enclosure
0158 UT   Sunset.  Opening telescope and beginning pointing check.
0202 UT   Resetting GCS due to MODS in calibration mode.
0206 UT   Difficulty acquiring 4-th magnitude pointing check star.
0214 UT   Finished first pointing check.  Moving to fainter star to check focus
0216 UT   Moving to pointing star near first object (Feige 34)
0220 UT   Unable to focus on pointing star.  Sky bkg too high?
0229 UT   Moving to acquire Feige 34
0234 UT   Lost guide star.  May be attempting to observe through cloud cover
0238 UT   "Parter CALIBRATION timeout"  Refresh MODS controller.
0252 UT   Slewing to SN 2012aw field
0400 UT   Slewing to SN 2007rt field.  Unable to maintain guiding
0412 UT   Slewing to SN 2005bf field.  Unable to locate guide star.  Clouds
0455 UT   Taking spectrum at SN 2005bf, but through clouds.  Data will be poor
0550 UT   Taking additional spectra to make up for the limited signal in the first two spectra
0628 UT   Moving telescop to Bootes field to begin EAGES_487588
0647 UT   Removed recommended rotation adjustment due to slight offsets between reference stars.  This did not appear to help.
0653 UT   Re-sending the preset command due to collimation error by AGw.  Failure.  Appears to have resulted from pupil crossing lenslets.
0657 UT   Sending AGw to on-axis position w/ bright star to check collimation
0710 UT   Sending telescope to SN 2012au field.  Unable to acquire guide star.
0730 UT   Attempting to acquire SN 2012au
0748 UT   Beginning observations of SN 2012au through variable clouds.  Guide star magnitude highly variable.
0750 UT   Lost running log entry due to wiki error.  Attempting to recover from separate file, but some log entries will be absent.
0844 UT   Sending telescope to HZ 44 field to take spectrophotometric standard
0849 UT   Beginning standard star spectra
0907 UT   Moving telescope to SN 2010mc reference object & beginning acquisition.
0910 UT   IIF Error.  Trying to send a manual preset to a nearby star.  
0912 UT   Failed to begin guiding for nearby star.
0920 UT   Resetting GCS to attempt to clear the problem.
0924 UT   GCS working again.  Collimating telescope & returning to SN 2010mc.
0939 UT   Beginning spectra of SN 2010mc
1000 UT   Sky is essentially clear with only light cirrus.
1047 UT   Slewing to Bootes field for EAGES_487588.
1105 UT   Beginning spectra of EAGES_487588
1107 UT   Position of guide star pupil image jumped again, but collimation appears to be okay so far.
1150 UT   Finished second integration for EAGES_487588.  Postponing additional integrations due to encroaching twilight. Beginning calibrations
1200 UT   Conducting spectral flats & arc lamp calibration
1257 UT   Finished MOS & longslit flats and some of the slitless flats.  Unable to finish all calibrations prior to handover.

David Atlee -- 2013 April 12
Topic revision: r4 - 12 Apr 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback